A spring of force constant k is cut into three equal parts. What is the force constant of each part ?

Grade:

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Each will have a force constant of 3k..

1/k=1/3k   + 1/3k  + 1/3k

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com

vikas askiitian expert
13 years ago

length of each part = L/3

 K inversly proportional to length so

K/K1 = L1/L

K1 = KL/L1 = KL/L/3 = 3K

each part will have force constant = 3k

Kushagra Madhukar
3 years ago
Dear student,
 
length of each part = L/3
SInce, K inversly proportional to length so
K/K* = L*/L

K* = KL/L* = KL/(L/3) = 3K

each part will have force constant = 3k

 

Thanks and regards,

Kushagra